OpenStudy (anonymous):

What is the volume of a sphere that has a diameter of 12 mm?

OpenStudy (anonymous):

V≈9.05×10-7m³

OpenStudy (anonymous):

@KeyriMelissa

OpenStudy (anonymous):

V=4/3πr3 V=4/3*3.14*6^3 V=904.78

OpenStudy (anonymous):

round up and its 9.05

OpenStudy (anonymous):

Good job @Night-Watcher

OpenStudy (anonymous):

@chaser71 @Night-Watcher thats not an answer choice

OpenStudy (anonymous):

Thanks Chaser! :)

OpenStudy (anonymous):

Can you put them?

OpenStudy (anonymous):

it has to be... if you correctly work this out then it should be 9.04 which rounds up to 9.05

OpenStudy (anonymous):

what are they? is one, 7238?

OpenStudy (anonymous):

@KeyriMelissa

OpenStudy (anonymous):

im working on it now and if you say yes then Ill show you how .

OpenStudy (anonymous):

A.288sqrt B.176sqrt C.324sqrt D.396sqrt @chaser71

OpenStudy (anonymous):

it needs to be in square root? not mm????

OpenStudy (anonymous):

@chaser71 i got its 288 but thank you